Three boxes contain lamp bulbs some of which are defective. The proportion of

defective in box A1; A2 and A3 are 1 2; 1 8 and 3 4 respectively. A box is selected at random and

a bulb is drawn from it. If the selected bulb is found to be defective, what is the probability that

box A1 was selected?


1
Expert's answer
2021-05-24T16:03:48-0400

Given that "A_1" ,"A_2" and "A_3" represent the three boxes.

Then "P(A_1" ")=P(A_2)=P(A_3)=" "\frac{1}{30}"


Let Q be the event of drawing the defective bulb.

Then "P\left(\frac{Q}{A_1}\right)""=P\left(drawing\:a\:defective\:bulb\:from\:A_1\right)=\frac{1}{2}"

"P\left(\frac{Q}{A_2}\right)=P\left(drawing\:a\:defective\:bulb\:from\:A_2\right)=\frac{1}{8}"

"P\left(\frac{Q}{A_3}\right)=P\left(drawing\:a\:defective\:bulb\:from\:A_3\right)=\frac{3}{4}"


The probability of selecting a defective bulb from "A_1" being given that it is defective , is "P\left(\frac{A_1}{Q}\right)."

Then "P\left(\frac{A_1}{Q}\right)= \frac{P\left(A_1\right)\times P\left(\frac{Q}{A_1}\right)}{\left[P\left(A_1\right)\times P\left(\frac{Q}{A_1}\right)\right]+\left[P\left(A_2\right)\times \:P\left(\frac{Q}{A_2}\right)\right]+\left[P\left(A_3\right)\times \:P\left(\frac{Q}{A_3}\right)\right]}"

"P\left(\frac{A_1}{Q}\right)=" "\frac{\frac{1}{3}\times \frac{1}{2}}{\left(\frac{1}{3}\times \:\frac{1}{2}\right)+\left(\frac{1}{3}\times \frac{1}{8}\right)+\left(\frac{1}{3}\times \frac{3}{4}\right)}"


"P\left(\frac{A_1}{Q}\right)= \frac{\frac{1}{2}}{\frac{1}{2}+\frac{1}{8}+\frac{3}{4}}=\frac{\frac{1}{2}}{\frac{4+1+6}{8}}=\frac{\frac{1}{2}}{\frac{11}{8}}"

"P\left(\frac{A_1}{Q}\right)=\frac{1}{2}\times \frac{8}{11}=\frac{4}{11}" .